OpenStudy (anonymous):

How can I solve this rational expressions: p^2 - 1/5p + 25 = p/5

DivineSolar (divinesolar):

You are looking for a real solution correct?

OpenStudy (anonymous):

Yes.

DivineSolar (divinesolar):

Alright. This will take some time :)!

OpenStudy (anonymous):

Okay.

DivineSolar (divinesolar):

Step 1 is to simplify both sides of the equation. Then we subtract 1/5p from both sides which looks like -> \[p^2 + \frac{ -1 }{ 5 }p + 25 - \frac{ 1 }{ 5 }p = \frac{ 1 }{ 5 }p - \frac{ 1 }{ 5 }p\]

DivineSolar (divinesolar):

Which turns into what?

OpenStudy (anonymous):

Could I factor the equation, or should I not?

DivineSolar (divinesolar):

Do not factor.

DivineSolar (divinesolar):

So this transfers to -> \[p^2 + \frac{ -2 }{ 5 }p + 25 = 0\]

DivineSolar (divinesolar):

See how i got that?

OpenStudy (amistre64):

my approach would be to multiply by 5 and subtract a p ...

DivineSolar (divinesolar):

Amistre lol.

DivineSolar (divinesolar):

We are trying to see if it has a real solution or not :) which i already know.

OpenStudy (amistre64):

:) youre doing fine then

DivineSolar (divinesolar):

@5jackyh do you see how i got that last answer?

DivineSolar (divinesolar):

?

OpenStudy (anonymous):

I'm working on it.

DivineSolar (divinesolar):

Mkay. Because for the next part we are using the quadratic formula where a = 1 , b = -0.4, c = 25

DivineSolar (divinesolar):

You there? :)?

DivineSolar (divinesolar):

Here is the work for quadratics, based upon this info what would your answer be @5jackyh ?

DivineSolar (divinesolar):

hmm?

OpenStudy (anonymous):

I got it.

DivineSolar (divinesolar):

What is it :)?

DivineSolar (divinesolar):

Jack?

DivineSolar (divinesolar):

Well basically it is so easy as i already gave it out with the quadratics, it would be no real solution.

OpenStudy (amistre64):

p^2 - 1/5p + 25 = p/5 5p^2 -p +125 = p 5p^2 -2p +125 = 0 a b c if b^2 - 4ac is negative, then no real roots :) 4 -4(5)(125) ... yeah, its definitely gonna be negative ....